Parkes observatory report

Operations

Operations in the last five months have been generally very smooth, with the sole exception of strong winds throughout much of December and January. This has become a very strong annual pattern in recent years.

In particular there has been no recurrence of the problems with the zenith drives experienced in September 2006. The time lost over the last five months is dominated entirely by high winds, at 5.2% of all time lost, with only 0.5% lost to faults.

The GASS (Galactic All-sky Survey) completed observations at Parkes in November, after a two-year observing program. We all look forward to seeing the final images.

Pilot observations for a new survey of continuum polarisation at a wavelength of 13 cm, dubbed S-PASS (S-Band Polarization All Sky Survey) commenced in January. This observing program will scan the sky with the 64-metre telescope at a slew rate of 15 degrees per minute in azimuth, the fastest scanning rate for a large survey that has been employed at Parkes for some time (or perhaps ever). The survey will certainly give the new azimuth gears a good workout when it commences in earnest.

Duplication of the entire Parkes 20-cm Galactic Pulsar Survey (P268) database is underway. It is expected that a complete copy of this approximately five-Terabyte database at Jodrell Bank will be available at Parkes later this year.

November-December shutdown

The 64-metre telescope was off the air for three weeks from 28 November 2006 for a long-planned refurbishment of the azimuth gearboxes. Essentially all moving parts in these gearboxes, most of them 45 years old, were replaced with new components. This work was successfully completed on time, and the new gearboxes are being closely monitored during their running-in. This work was carried out with assistance from Brian Wilcockson and Steve Broadhurst from Marsfield, Tim Wilson from Narrabri and of course Jon Crocker from Parkes.

A second major work in the shutdown was the removal of the 20-cm multibeam package for the second phase of its refurbishment. The removal went very smoothly and the receiver is currently disassembled in the receiver laboratory as the new low noise amplifiers (LNAs) are installed and other remedial works carried out. A fault was discovered in several of the ortho mode transducers (OMTs) which is strongly suspected of causing the microphonic instabilities observed in several of the channels over recent years. There is some confidence that these defects, present since construction, have now been completely eliminated. The receiver is due to return to service in April or May 2007.

Some significant changes to the master equatorial (ME) drive system were implemented, including replacement of old (and mostly undocumented) drive electronics, and replacement of the long-serving master equatorial computer (PDP 11/20) with another 11/20 chassis of a type for which we have many more spare parts available. New software installed by Andrew Hunt has improved the tracking characteristics of the ME at the arcsecond level. Modifications to utilise the 13th bit of precision (8192 counts/turn) are mostly complete, and will be finalised at the next significant shutdown.

Other works carried out in the shutdown included installation of a new sliding receiver hatch on the focus cabin roof and installation of new airconditioning for the compressor room, both of which will improve the efficiency and reliability of operations. The first floor control room and stairs have been recarpeted and the "Olympic Village" has at last received its roof, giving its inhabitants rain- and sun-free access to the administration building.

Further upgrading of the mains switchgear was carried out to pave the way for full mains-synchronisation of the observatory generator supply. This project, to be completed later in the year, will allow seamless cutover from mains to generator — a great operational improvement.

Numerous other improvements around the site were also carried out during the shutdown period, all under the capable supervision of Barry Turner, our site-services manager.

Computing

Advantage was taken of the shutdown to install a new core switch (Cisco 6503) for the site, with work undertaken by Shaun Amy and Kan Tam. The new switch provides for an increase in the number of 1Gb/s connections to machines on-site, to take better advantage of the new broadband observatory links.

The shutdown was also used to repair optic fibres that had failed some months previously from rodent damage. This task was far from trivial, fully occupying a large team of cabling contractors for two days. The damage highlighted some inadequacies in the original cabling upgrade in 2004, which were unfailingly pinpointed by local wildlife. Brett Preisig was as usual on hand to ensure the operation ran smoothly and to specifications. Rodent counter-measures have been escalated on all fronts to prevent similar problems elsewhere on site.

Quarters

The final phase of the observers' Quarters upgrade was completed in November. This last stage comprised painting and repolishing of the timber in the enlarged dining area. The new facilities are a great credit to those involved in this project, all of whom are very glad to see it completed!

John Reynolds
Officer-in-Charge, Parkes Observatory
(John.Reynolds@csiro.au)

News
Public
Please provide your user name and password | Australia Telescope National Facility
 

Please provide your user name and password

Enter your CSIRO NEXUS username
Enter your NEXUS password
Access denied | Australia Telescope National Facility
 

Access denied

You are not authorized to access this page.
Access denied | Australia Telescope National Facility
 

Access denied

You are not authorized to access this page.